OpenStudy (littlenugget):

Given the equation of the line, 2x-3y=7, find each of the following: a) an equation of the line through (2,-3) parallel to the given line. b) an equation of the line through (-3,2) perpendicular to the given line.

OpenStudy (john_es):

(a) First, I would express the equation given in this form y=mx+n, \[y=\frac{2}{3}x-\frac{7}{3}\]As the equation must be parallel, then \[y=\frac{2}{3}x+n\] And now you force it to pass by (2,-3) \[-3=\frac{4}{3}+n\Rightarrow n=-13/3\] (b) For the second problem, follow the same path, but now the equation of the perpendicular will be y=-x/m+n, so, \[y=-\frac{3}{2}x+n\]And force it to pass by (-3,2), \[2=-9/2+n\Rightarrow n=13/2\]

OpenStudy (littlenugget):

now did you get the n?

OpenStudy (littlenugget):

@John_ES

OpenStudy (john_es):

The n in the first case was -13/3, and in the second 13/2, if I'm not mistaken.

OpenStudy (littlenugget):

ohhhh i see how you got it :) thanks!!! @John_ES